OpenStudy (anonymous):

54^x2 - 6 Factor each completely.

OpenStudy (anonymous):

Well there's no x in the second term so there's nothing to worry about there. Can you find a common factor for the coefficients (54 and 6)

OpenStudy (anonymous):

It would be 3 right?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

6* lol i'm tired

OpenStudy (anonymous):

Yeah 6. And then i would do what with the 6?

OpenStudy (kuoministers):

You take 6 out of the equation. 6 (9x^2 - 1) So can you see if you can factorise this any more? hint: A - B = (a+b)(a-b)

OpenStudy (anonymous):

i dont understand what youre suppposed to do next, can you explain please

OpenStudy (anonymous):

Actually i got it :)

OpenStudy (anonymous):

It would be 6(3x+1) (3x-1)

OpenStudy (kuoministers):

6(3x-1)(3x+1) This is because of the rule (a+b)(a-b) = A - B Ok :)

OpenStudy (anonymous):

okay well thank you! :)